OpenStudy (anonymous):

Which of the following sets could be the sides of a right triangle? a. {5, 8, 12}

OpenStudy (anonymous):

b.{2, 3,\[\sqrt{10}\]} c.{3, 5,\[\sqrt{34}\]}

OpenStudy (anonymous):

The valid set must comply with Pythagoras Theorem which means that in the set there must be a number whose square is the sum of the squares of the other two. If I take the first option {5, 8, 12} and calculate the squares, I get {25,64,144} and there is no way to make any of hese numbers be the sum of the other two: 25+64=89<>144 25+144=169<>64 64+144=208<>25 Then, option a) is not valid Play with the other choices and get back to me with your proposed solution

OpenStudy (anonymous):

theres b. {2,3, \[\sqrt{10}\]} and c. {3,5,\[\sqrt{34}\]}

OpenStudy (anonymous):

I know, but I am not going to give you the answer. I have given you the clue to find out yourself. Repeat what I hhave done for the other two options

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

idk.. can you tell me how you did it step by step?

OpenStudy (anonymous):

Take option b) and give me the sqaure of each one of the numbers